### v2

# DiScoFormer: Plug-In Density and Score Estimation with Transformers

Abstract
Estimating probability density and its score from samples remains a core problem in generative modeling, Bayesian inference, and kinetic theory. Existing methods are bifurcated: classical kernel density estimators (KDE) generalize across distributions but suffer from the curse of dimensionality, while modern neural score models achieve high precision but require retraining for every target distribution. We introduce DiScoFormer (Density and Score Transformer), a “train-once, infer-anywhere” equivariant Transformer that maps i.i.d. samples to both density values and score vectors, generalizing across distributions and sample sizes. Analytically, we prove that self-attention can recover normalized KDE, establishing it as a functional generalization of kernel methods; empirically, individual attention heads learn multi-scale, kernel-like behaviors. The model converges faster and achieves higher precision than KDE for density estimation, and provides a high-fidelity plug-in score oracle for score-debiased KDE, Fisher information computation, and Fokker-Planck-type PDEs.
density estimation,score estimation,transformers,kernel methods,equivariant networks,nonparametric statistics
1 Introduction
Estimating a probability density $f$ and its corresponding score function $\nabla\log f$ from i.i.d. samples is a foundational problem in statistical inference, underpinning tasks in generative modeling, stochastic control, and information theory. While classical non-parametric methods, such as Kernel Density Estimation (KDE), offer strong theoretical guarantees and respect fundamental symmetries, they are hampered by a rigid bias-variance trade-off and a prohibitive “curse of dimensionality” as dimension $d$ increases. Conversely, modern neural score-matching models achieve high accuracy in high-dimensional settings but are typically transductive: they must be retrained for every new target distribution, precluding their use as general-purpose, “off-the-shelf” estimators.
We propose bridging this gap by framing density and score estimation as a sequence-to-operator (Sutskever et al., 2014) learning task. Given an i.i.d. sample $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$, we seek to learn a single model that maps the entire sample $X$ to its underlying density and score functions. Specifically, we define an operator $T$ for density and $S$ for scores:
$\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\xrightarrow{\;T\;}\begin{pmatrix}f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] f(x_{n})\end{pmatrix},\qquad\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\xrightarrow{\;S\;}\begin{pmatrix}\nabla\log f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla\log f(x_{n})\end{pmatrix}$. (1)
A critical requirement for such an operator is that it must respect the inherent symmetries of the data. Specifically, the estimates should be permutation-equivariant with respect to the sample index and affine-equivariant with respect to the coordinate space. For a permutation matrix $P$, invertible matrix $A$, and shift $\mu$, these symmetries are defined as:
$T(PXA+\mathbf{1}\mu^{\top})$ $=P\,|\det A|^{-1}\,T(X),$ (2)
$S(PXA+\mathbf{1}\mu^{\top})$ $=P\,A^{-1}\,S(X),$ (3)
Building on these principles, we introduce DiScoFormer – Density and Score Transformer. By treating the sample $X$ as a sequence, the Transformer architecture provides permutation equivariance by construction. We achieve affine equivariance through a specialized whitening mechanism that normalizes inputs up to orthogonal transformations, combined with data augmentation to ensure rotation invariance. Unlike traditional models, our architecture utilizes cross-attention, allowing for the evaluation of $f$ and $\nabla\log f$ at arbitrary query points without retraining.
Furthermore, we provide a theoretical bridge between modern attention mechanisms and classical non-parametrics. We demonstrate analytically that under specific parameterizations, self-attention weights recover normalized KDE weights. This expressivity result positions the Transformer not as a black box, but as a principled, data-adaptive generalization of kernel methods that can dynamically learn multiscale smoothing behaviors from data.
Empirically, the model outperforms KDE and score-debiased KDE (Epstein et al., 2025) across a range of dimensions and sample sizes on both in-distribution and out-of-distribution test cases. Trained on Gaussian Mixture Model (GMM) data, the model generalizes to GMMs with more modes and non-Gaussian targets such as the Laplace distribution and demonstrates favorable scaling in both $n$ and $d$. By amortizing the estimation cost, our model provides high-fidelity plug-in oracles for downstream tasks, including entropy estimation, Fisher information calculation, and solving Fokker-Planck-type partial differential equations (PDEs).
In summary, our contributions are:
  • DiScoFormer, a universal Transformer model for one-shot estimation of density and score functions from i.i.d. samples.
  • A proof and empirical evidence establishing self-attention as a data-adaptive generalization of normalized kernel density estimation.
  • Non-parametric density and score estimators that surpass classical KDE methods in accuracy across sample sizes and dimensions.
  • Applications to Fisher information, differential entropy, and Fokker-Planck-type PDEs.
2 Related Work
Our work lies at the intersection of nonparametric density estimation, score-based methods, permutation-invariant neural architectures, and operator learning for probability measures.
Kernel density estimation.
Classical nonparametric methods such as Parzen windows (Parzen, 1962) and Silverman’s rule-of-thumb KDE (Silverman, 1986; Wand and Jones, 1994) estimate densities via local kernel averaging. While theoretically grounded, these methods suffer from a rigid bias–variance trade-off and scale poorly with dimension (Scott, 2015). Adaptive bandwidth methods (Abramson, 1982) and fast algorithms such as the Fast Gauss Transform (Greengard and Strain, 1991) partially alleviate computational costs but do not resolve the fundamental curse of dimensionality. Score-Debiased KDE (Epstein et al., 2025) demonstrates that access to a score oracle enables bias correction, motivating our approach: we provide precisely this missing oracle as a reusable, distribution-agnostic component.
Other nonparametric density estimators.
$k$-nearest-neighbor methods (Loftsgaarden and Quesenberry, 1965), orthogonal series estimators (Efromovich, 2010; Scott, 1985), and normalizing flows (Rezende and Mohamed, 2015; Dinh et al., 2017; Papamakarios et al., 2017) offer alternatives to KDE but either lack direct score estimates or require per-distribution retraining. Our model provides both density and score from a single pretrained network.
Score matching and score-based generative models.
Score matching (Hyvärinen, 2005) and its denoising variant (Vincent, 2011) estimate unnormalized model scores without computing partition functions. Sliced score matching (Song et al., 2019) scales these ideas to higher dimensions. Score-based generative models (Song and Ermon, 2019; Ho et al., 2020; Song et al., 2021b, a) learn scores along diffusion processes for sampling, with fast solvers (Lu et al., 2022; Karras et al., 2022) and convergence guarantees (Chen et al., 2023, 2024). These methods train on a single target density; in contrast, our model learns a universal score operator that generalizes across distributions without retraining.
Permutation-invariant architectures.
DeepSets (Zaheer et al., 2017) and Set Transformers (Lee et al., 2019) provide foundational architectures for exchangeable data, while equivariant graph networks (Maron et al., 2019) extend these ideas to richer symmetry groups. Neural Processes (Garnelo et al., 2018; Edwards and Storkey, 2017) and Perceiver IO (Jaegle et al., 2022) learn distributions over functions from context sets. Recent work shows that permutation-equivariant transformers can universally approximate mean-field dynamics (Biswal et al., 2025). Our architecture exploits the same symmetry but outputs statistical quantities—density and score—while additionally enforcing affine equivariance.
Attention as kernel smoothing.
Formal connections between softmax attention and kernel methods have been developed through linear-attention reformulations (Katharopoulos et al., 2020), random-feature approximations (Choromanski et al., 2021), Nyström methods (Xiong et al., 2021), explicit Gaussian-kernel replacements (Chen et al., 2021; Lu et al., 2021a), and Nadaraya–Watson regression interpretations (Zhang et al., 2023). Our Proposition 3.2 builds on this lineage by showing that self-attention weights can exactly reproduce normalized Gaussian KDE weights.
Neural operators.
Neural Operators (Kovachki et al., 2023), DeepONet (Lu et al., 2021b), and Fourier Neural Operators (Li et al., 2021) learn mappings between infinite-dimensional function spaces. The Score Neural Operator (Liao et al., 2024) extends this framework to probability distributions via RKHS embeddings. In contrast, our transformer acts directly on raw i.i.d. samples without kernel embeddings, yielding a simpler operator that generalizes across densities and dimensions.
Particle methods and Fokker–Planck solvers.
Interacting particle methods for Fokker–Planck equations (Carrillo et al., 2019; Maoutsa et al., 2020) and score-based transport modeling (Boffi and Vanden-Eijnden, 2023; Lu et al., 2024; Ilin et al., 2025a) require accurate score estimates at each timestep. Stein Variational Gradient Descent (Liu and Wang, 2016; Liu, 2017; Duncan et al., 2023) similarly relies on kernel-based score approximations. Our model provides a fast, pretrained oracle that plugs directly into these solvers, avoiding per-distribution retraining.
3 Methodology
Figure 1: Attention visualization of the transformer model. The top panel shows the average attention scores in layer 0 with respect to the chosen point x. The heatmaps show the attention matrix and the normalized KDE matrix $D_{i,j}\propto e^{-\|x_{i}-x_{j}\|^{2}_{2}}$. The scatter plots show very high agreement between attention scores and $D$. The bottom four panels visualize the attention scores of individual heads, demonstrating emergent head specialization.
In this section we describe the core ideas of the proposed method: the symmetry-aware architecture, the relationship to classical kernel density estimation, the training pipeline, and implementation.
3.1 Equivariance
Table 1: Relative MSE of affine equivariance error, averaged over 50 trials.
Transform Rel. MSE
Permutation 0
Translation 0
Isotropic scaling 0
Anisotropic scaling 0
Rotation $5\times 10^{-4}$
Full affine $1\times 10^{-4}$
First, Proposition 3.1 establishes the equivariance properties of density and score estimation. The proof is in the Appendix.
Proposition 3.1 (Permutation and affine equivariance of density and score evaluation).
Let $f$ be a differentiable density, and $X=(x_{1},\dots,x_{n})^{T}$ be its iid sample. Define
$T(X):=\begin{pmatrix}f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] f(x_{n})\end{pmatrix},\qquad S(X):=\begin{pmatrix}\nabla\log f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla\log f(x_{n})\end{pmatrix}$.
Let $P\in\mathbb{R}^{n\times n}$ be a permutation matrix, $A\in\mathbb{R}^{d\times d}$ be invertible, $\mu\in\mathbb{R}^{d}$, and $\mathbf{1}\in\mathbb{R}^{n}$ be the vector of ones. Then
$T\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,|\det A|^{-1}\,T(X),$ (4)
$S\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,A^{-1}\,S(X).$ (5)
To capture these symmetries, we use the Transformer architecture without positional encodings, combined with an affine normalization layer (Figure 2). The whitening step yields equivariance up to rigid rotations $O(d)$; training on randomly oriented GMMs encourages the network to learn approximate rotation invariance, closing the remaining gap in practice. Table 1 confirms this empirically.
m = X.mean(0, keepdim=True)
Xc, Yc = X - m, Y - m
S = Xc.T @ Xc + eps * torch.eye(d)
A = matrix_sqrt_inv(S)
Xw, Yw = Xc @ A, Yc @ A
dens, score = self._core(Xw, Yw)
return dens * torch.det(A), score @ A.T
Figure 2: The forward pass implementing affine equivariance.
3.2 Attention and KDE
Proposition 3.2 demonstrates that a single attention head can compute the classical KDE, showing that the attention mechanism has the capacity to implement kernel-based smoothing. While the proposition assumes $L^{2}$-normalized inputs, in practice the learned $Q$ and $K$ projections adapt to the actual input statistics after whitening.
Proposition 3.2 (Attention can compute KDE).
Let $X\in\mathbb{R}^{n\times d}$ be the normalized data matrix with $\|x_{i}\|=1$ for all $i$. Consider a single self-attention head
$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(QK^{\top})\,V$
with
$Q=K=\tfrac{1}{h}X,\qquad V=I_{n},$
for some bandwidth parameter $h>0$. Then
$\mathrm{Attention}(Q,K,V)]_{ij}=\frac{\exp\!\left(-\frac{\|x_{i}-x_{j}\|^{2}}{2h^{2}}\right)}{\sum_{k=1}^{n}\exp\!\left(-\frac{\|x_{i}-x_{k}\|^{2}}{2h^{2}}\right)},$
i.e. the attention matrix coincides with the normalized Gaussian kernel of bandwidth $h$.
While the above Proposition establishes that the Attention mechanism has the capacity to approximate KDE well, we must verify whether the model learns to do it practice. To empirically verify the relationship between the KDE weights and attention scores, we visualize attention in Figure 1. As expected, the Transformer learns to attend to nearby points, with one attention head attending over far away samples instead. The attention mechanism effectively learns a kernel-like behavior, with emergent head specialization. The learned features allow for much better scaling both in the sample size and dimension, compared to KDE. In section B of the Appendix we further visualize the attention scores of individual heads, and find an emergent behavior of head specialization. Specifically, head 1 specialized to look at far-away points, whereas heads 0, 2, and 5 specialized to look at close- and mid-range interactions. Finally, heads 3, 4, 6, and 7 specialized to look in specific directions. This emergent behavior further links the multi-head Transformer to kernel-based methods, but with flexible multi-scale kernels.
3.3 Training Data
We train the Transformer to approximate the density and score of Gaussian Mixture Models (GMMs) from i.i.d. samples. GMMs are chosen because they form a dense family in the space of continuously differentiable densities, and both their densities and scores admit closed-form expressions. Training samples are generated on the fly (Algorithm 1). The objective combines the mean squared errors (MSE) of the predicted density and score through a convex weighting:
$\mathcal{L}_{T}$ $=\frac{1}{n}\|T(X,Y)-f_{X}(Y)\|_{2}^{2},$ (6)
$\mathcal{L}_{S}$ $=\frac{1}{n}\|S(X,Y)-\nabla\log f_{X}(Y)\|_{2}^{2}$ (7)
$\mathcal{L}$ $=\alpha\mathcal{L}_{T}+(1-\alpha)\mathcal{L}_{S}.$ (8)
Algorithm 1 GMM DataLoader
Input: batch size $B$, dimension $d$, sample sizes $n_{x},n_{y}$, components $k_{\min},k_{\max}]$
repeat
  Sample $k\in\{k_{\min},\dots,k_{\max}\}$
  for $b=1$ to $B$ do
   Sample two random GMMs with $k$ components each
   Sample $X_{b}$ from the first, $Y_{b}$ from the second
   Compute $f_{X_{b}}(y)$ and $\nabla\log f_{X_{b}}(y)$ for $y\in Y_{b}$
  end for
  Output: $\big(X,\,Y,\,f_{X}(Y),\,\nabla\log f_{X}(Y)\big)$
until training stops
3.4 Model Variants and Implementation
We explore several architectural variants aimed at extending density and score estimation beyond the observed samples $X$, enabling evaluation at arbitrary query points $Y$. To achieve this, we introduce a cross-attention mechanism that processes two complementary inputs: a context tensor encoding the observed samples and a query tensor specifying the target locations. Through this interaction, the model learns to relate queries to the statistical structure of the observed data, producing accurate estimates of both $f_{X}(y)$ and $\nabla\log f_{X}(y)$ for any $y\in Y$. The cross-attention layer thus serves as a learned nonparametric smoother, effectively generalizing density and score estimation across the input domain rather than being limited to the training samples.
Recognizing that density and score estimation are mathematically coupled – linked through the gradient of $\log f$ – we employ a unified architecture with a shared backbone and two specialized output heads. This joint formulation encourages mutual consistency between the two quantities and allows the model to leverage shared geometric and statistical features during training, improving both accuracy and sample efficiency.
4 Experiments
Figure 3: Score estimation comparison between Silverman KDE (Silverman, 1986) and our transformer model. The transformer is more accurate, especially in the sparse regions. We plot the negated score for easier viewing.
In this section, we empirically investigate:
  1. 1.
    comparison with KDE and SD-KDE (Epstein et al., 2025),
  2. 2.
    scaling in sample size $n$ and dimension $d$,
  3. 3.
    out-of-distribution generalization,
  4. 4.
    inference-time scaling in $n$ and $d$.
We also show that the learned model can serve as a score oracle in SD-KDE, as well as in deterministic SBTM-like solvers for Fokker-Planck-type equations such as the Landau equation.
While there are many techniques for learning the score $\nabla\log f$ such as ISSM (Song et al., 2021a, b) and DDPM (Lu et al., 2022), they learn the score of a specific density undergoing a forward diffusion-like stochastic process. Since our model is density-agnostic, and can be used out-of-the-box on a sample of any density, the most comparable method is KDE and its variants. See the Appendix for a comparison between the score matching loss and our model.
All experiments are performed on a single 48GB L40S GPU. The architecture is a permutation- and affine-equivariant Transformer. It consists of a linear embedding, four Transformer encoder layers (128-dimensional hidden size, 8 heads, GELU activation, pre-normalization), and a linear projection to the output dimension. No positional encodings are used, preserving permutation-equivariance. The model has about $800,000$ parameters. We use batch size $32$, sample size $n=2048$, dropout $0.1$, and draw GMMs with 1–10 modes with means in $3,3]^{d}$ and diagonal covariances in $0,1]^{d}$. A runtime comparison with KDE is provided in Appendix D.
4.1 Score estimation
Figure 4: MSE of score estimation on the Laplace distribution. Test-time training (TTT) improves the out-of-distribution generalization.
To compare our model to KDE, we first sample a 2d Gaussian, visualize the estimated scores using a quiver plot and compute the MSE score loss. Figure 3 demonstrates that even in dimension 2 and a KDE-favorable example of a unimodal Gaussian, our model performs better. The KDE score approximation is computed using
$k(x_{i},x_{j})$ $=\exp\!\left(-\frac{1}{2}\sum_{\ell=1}^{d}\frac{(x_{i}^{\ell}-x_{j}^{\ell})^{2}}{h_{\ell}^{2}}\right),$ (9)
$\widehat{s}(x_{i})$ $=\frac{1}{h^{2}}\odot\left(\frac{\sum_{j}k(x_{i},x_{j})x_{j}}{\sum_{j}k(x_{i},x_{j})}-x_{i}\right),$ (10)
where $h_{\ell}=\sigma_{\ell}n^{-1/(d+4)}$ is Silverman’s rule (Silverman, 1986) applied per coordinate with $\sigma_{\ell}$ the sample standard deviation along dimension $\ell$, and $\odot$ denotes element-wise multiplication.
For a quantitative comparison of score estimation against the KDE, we compute the Mean Squared Error (MSE) for dimensions $1$ and $10$ across several magnitudes of sample sizes $n$. See the left panel of Figure 5. The transformer model achieves much better accuracy than KDE even in dimension 1 and especially in dimension 10.
To judge the out-of-distribution generalization of our model, we compare the MSE loss on GMMs with 1-19 modes, while the training data consists of GMMs with 1-10 modes. The right panel of Figure 5 shows that the MSE loss is monotone and stable in the number of modes. This indicates good generalization. Additionally, we evaluate on two non-Gaussian distributions: the Laplace distribution, which has a sharper peak and heavier tails than a Gaussian, and the Student-$t$ distribution ($\nu=3$), which has polynomial rather than exponential tail decay. Despite being trained only on GMMs, the model estimates the score successfully on both, as shown in Tables 2 and 3. Additionally, we can further improve the out-of-distribution performance using test-time training (TTT) (Sun et al., 2020; Gandelsman et al., 2022), using the consistency loss
$\mathcal{L}_{\text{con}}=\frac{1}{n}\sum_{i=1}^{n}\left\|S(X)_{i}-\nabla_{x_{i}}\log T(X)_{i}\right\|_{2}^{2},$ (11)
where $T(X)_{i}$ is the predicted density and $\nabla\log T(X)$ is computed via automatic differentiation. Figure 4 and Table 3 show that as few as 4 TTT steps are sufficient to improve performance on both distributions.
Table 2: MSE of score estimation on the 2D Laplace distribution.
$n$ KDE Transformer
512 0.3810 0.3598
1024 0.3305 0.2992
2048 0.2990 0.2756
4096 0.2650 0.2597
Table 3: MSE of score estimation on the 2D Student-$t$ distribution ($\nu=3$).
$n$ KDE No TTT TTT 4 TTT 6 TTT 8
128 0.1515 0.1980 0.1676 0.1569 0.1450
256 0.1206 0.1119 0.0956 0.0908 0.0897
512 0.0916 0.0574 0.0488 0.0485 0.0514
1024 0.0812 0.1023 0.0806 0.0768 0.0765
Figure 5: Left: MSE of score estimation in dimensions 1 and 10 on a 3-modal GMM using Transformer and KDE. The Transformer has excellent scaling in both dimension $d$ and the sample size $n$. Right: MSE of score estimation using the Transformer on GMMs with different numbers of modes. Despite being trained only on GMMs with 1-10 modes, and $n=2048$, the model exhibits excellent generalization.
4.2 Relative Fisher Information
The Relative Fisher Information (RFI) measures how two densities differ in their local geometry, comparing their score fields rather than their probability mass. Unlike the KL divergence, which captures global differences, RFI focuses on the mismatch between gradients of the log-densities, offering a more sensitive criterion for evaluating score-based models and generative estimators.
Figure 6: Computing relative Fisher information. Our model predicts $\nabla\log g(x_{i})$ at query points $x_{i}$ via cross-attention with the samples $y_{i}\sim g$. These predicted scores can be used to estimate the relative Fisher information.
Formally, for two densities $f$ and $g$ on $\mathbb{R}^{d}$, the relative Fisher information of $g$ with respect to $f$ is defined as
$\mathrm{I}(g\|f)=\mathbb{E}_{x\sim f}\!\left[\,\|\nabla_{x}\log g(x)-\nabla_{x}\log f(x)\|_{2}^{2}\,\right],$
and can be approximated by the Monte Carlo estimate on samples $x_{i}\sim f$ with
$\mathrm{I}(g\|f)\approx\frac{1}{n}\sum_{i=1}^{n}\|\nabla\log g(x_{i})-\nabla\log f(x_{i})\|_{2}^{2}.$
Because our model outputs the scores $\nabla\log g(x_{i})$ using samples $y_{i}$ from $g$, we can compute relative Fisher information between two distributions using their samples, as visualized in Figure 6.
4.3 Density Estimation via SD-KDE
Figure 7: Qualitative comparison: True density, Silverman KDE, Emp-SD-KDE, and SD-KDE with our learned score. Learned-score SD-KDE is visibly less biased.
Figure 8: 1D bimodal GMM ($n=2048$): SD-KDE with the learned score best approximates the density by MSE.
Figure 9: MSE of density estimation in 1D (left) and 2D (right). SD-KDE with our learned score and the Transformer model show the best scaling. The Transformer was trained only at $n=2048$.
Knowledge of the score $\nabla\log f$ can be exploited by score-debiased KDE (SD-KDE) (Epstein et al., 2025) to reduce the bias from $O(h^{2})$ to $O(h^{4})$. SD-KDE sharpens samples by moving them along the score before applying KDE:
$X\mapsto X+\frac{h^{2}}{2}\,\nabla\log f(X),$
which counteracts KDE’s smoothing bias. In practice the true score is unavailable. We therefore consider five methods: (i) plain KDE with Silverman bandwidth (Silverman, 1986); (ii) Empirical SD-KDE (Emp-SD-KDE), which uses the Silverman KDE to approximate the score; (iii) SD-KDE using our Transformer’s predicted score; (iv) SD-KDE using the autograd score of a Transformer-learned log-density; and (v) the Transformer’s direct density/log-density output.
  1. 1.
    KDE (Silverman) (Silverman, 1986).
  2. 2.
    Emp-SD-KDE: SD-KDE with the score derived from Silverman KDE, following (Epstein et al., 2025).
  3. 3.
    SD-KDE (Learned Score): SD-KDE with the Transformer-predicted score.
  4. 4.
    SD-KDE (Autograd Score): SD-KDE with the score from autograd of the Transformer’s learned log-density.
  5. 5.
    Transformer Density: direct density/log-density from our model.
Figure 7 shows that incorporating a learned score substantially improves density recovery over plain KDE; Emp-SD-KDE reduces bias, while SD-KDE with our predicted score yields the most accurate estimate. In 1D bimodal GMMs (Figure 8), SD-KDE with the learned score best matches the truth by MSE. Our Transformer-based approaches exhibit superior accuracy and scaling. The direct density approach performs best at small $n$ but, as expected, scaling with $n$ favors score-based methods; learning the score is an easier objective than learning the density itself due to the difficulty of learning the normalization constant $Z=\int_{\mathbb{R}^{d}}f$.
4.4 Estimation of entropy and Fisher information
Differential entropy $H(f)=-\mathbb{E}_{f}\log f$ and its gradient, Fisher information $I(f)=\mathbb{E}_{f}\|\nabla\log f\|_{2}^{2}$, are fundamental statistical quantities that characterize the spread of the distribution. We use the discretizations
$H(f)$ $\approx\frac{1}{n}\sum_{i=1}^{n}\log f(x_{i})\approx\frac{1}{n}\sum_{i=1}^{n}\log T(X)_{i}$ (12)
$I(f)$ $\approx\frac{1}{n}\sum_{i=1}^{n}\|\nabla\log f(x_{i})\|^{2}\approx\frac{1}{n}\sum_{i=1}^{n}\|S(X)_{i}\|^{2}$ (13)
to estimate $H(f)$ and $I(f)$ using only an iid sample of $f$. In dimension 1 we sample a GMM with three modes, numerically compute its entropy and Fisher information using expensive grid integration and the oracle knowledge of its density and score, and compute the estimation error. In dimension 10 we cannot afford to do grid estimation, so we sample a random Gaussian with randomized mean and covariance, and use the known analytical formulas. We use the debiased KDE method of (Epstein et al., 2025) here as well. Figure 10 demonstrates that already in dimension 1 both the score model $S$ and the density model $T$ outperform Silverman KDE. In dimension 10 the learned density performs much better for entropy estimation and the learned score performs much better for Fisher information estimation.
Figure 10: Comparison between the transformer model (learned) and Silverman KDE and score-debiased KDE for estimation of differential entropy $H(f)$ and Fisher Information $I(f)$. Transformer’s MSE is lower than that of the KDE approximation, even in dimension 1.
4.5 Plasma simulation
In physics, simulating plasma – hot ionized gas – is used in nuclear fusion and rocket propulsion. It is often done by solving the Landau equation, a Fokker-Planck-type kinetic PDE. In 2020 (Carrillo et al., 2020) recognized that the Landau equation can be rewritten in the form of a transport equation with velocity
$v(x)=-\int_{\mathbb{R}^{3}}A(x-y)(\nabla\log f(x)-\nabla\log f(y))f(y)dy$
and solved numerically with a particle method by approximating the score via a particular KDE variant, called the blob method. On the other hand, in machine learning Boffi and Vanden-Eijnden (2023) proposed a NN-based method of solving any Fokker-Planck equation called Score Based Transport Modeling, while Ilin et al. (2025a, b) adapted it to the Landau equation and the problem of sampling. The core idea is to continually train the NN $s_{\theta}$ on the score matching loss $\mathbb{E}_{f}\left[\|s_{\theta}(x)\|^{2}+2\nabla_{x}\cdot s_{\theta}(x)\right]$, which is equivalent to training on the MSE loss $\mathbb{E}_{f}\left[\|s_{\theta}-\nabla\log f\|^{2}\right]$. While Boffi and Vanden-Eijnden (2023); Ilin et al. (2025a); Ilin (2025) proved theoretically and demonstrated empirically that this online training effectively bounds the error, using SBTM can be slow due to the on-the-fly NN training. We reproduce the experiments 5.3 and 5.4 from (Ilin et al., 2025a) using the learned score. Unlike SBTM, we only use the pre-trained model and do not perform any training throughout the simulation. Figure 11 demonstrates that the Transformer model is able to match the analytic covariance well, while the KDE-based solver struggles.
Figure 11: Comparison between the trained Transformer model and Silverman KDE on the task of numerically solving the homogeneous Landau equation. We plot the first entry of the covariance matrix $\Sigma_{1,1}(t)$ of the numerical solutions and ground truth, when known. The left two panels use Maxwell collisions, while the last panel shows Coulomb collisions. The Transformer outperforms KDE, and is comparable in quality to SBTM in (Ilin et al., 2025a).
5 Conclusion
We introduced DiScoFormer, a permutation- and affine-equivariant Transformer that jointly estimates density and score from i.i.d. samples in a single forward pass. We proved that self-attention recovers normalized KDE and confirmed empirically that individual heads learn multi-scale kernel-like behaviors. DiScoFormer outperforms KDE across dimensions and sample sizes and serves as a plug-in oracle for score-debiased KDE, entropy and Fisher information estimation, and Fokker-Planck-type PDEs.
Appendix A Proofs
Proposition A.1 (Permutation and affine equivariance of density and score evaluation).
Let $f:\mathbb{R}^{d}\to(0,\infty)$ be a density, and for
$X=\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\in\mathbb{R}^{n\times d}$
define
$T(X):=\begin{pmatrix}f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] f(x_{n})\end{pmatrix},\qquad S(X):=\begin{pmatrix}\nabla\log f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla\log f(x_{n})\end{pmatrix}$.
Let $P\in\mathbb{R}^{n\times n}$ be a permutation matrix, $A\in\mathbb{R}^{d\times d}$ be invertible, $\mu\in\mathbb{R}^{d}$, and $\mathbf{1}\in\mathbb{R}^{n}$ be the vector of ones. Then
$T\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,|\det A|^{-1}\,T(X),$ (14)
$S\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,A^{-1}\,S(X).$ (15)
Proof.
Write $X=(x_{1}^{\top},\dots,x_{n}^{\top})^{\top}$ and let $\sigma$ be the permutation of $\{1,\dots,n\}$ corresponding to $P$, i.e. $(PX)_{i\cdot}=x_{\sigma(i)}^{\top}$. Define
$Y:=PXA+\mathbf{1}\mu^{\top}\in\mathbb{R}^{n\times d}.$
Then the $i$-th row of $Y$ is
[…]
Consider the density obtained from $f$ by the affine change of variables $y=A^{\top}x+\mu$:
$f^{(\mu,A)}(y):=|\det A|^{-1}f\big((A^{\top})^{-1}(y-\mu)\big).$
By construction,
$f^{(\mu,A)}(y_{i})=|\det A|^{-1}f\big(x_{\sigma(i)}\big).$
Hence, by the definition of $T$,
$T(Y)=\begin{pmatrix}f^{(\mu,A)}(y_{1})\\[-2.0pt] \vdots\\[-2.0pt] f^{(\mu,A)}(y_{n})\end{pmatrix}=\begin{pmatrix}|\det A|^{-1}f(x_{\sigma(1)})\\[-2.0pt] \vdots\\[-2.0pt] |\det A|^{-1}f(x_{\sigma(n)})\end{pmatrix}=P\,|\det A|^{-1}\,T(X),$
which proves (14).
For the score, differentiate $\log f^{(\mu,A)}$:
$\log f^{(\mu,A)}(y)=-\log|\det A|+\log f\big((A^{\top})^{-1}(y-\mu)\big).$
By the chain rule,
$\nabla_{y}\log f^{(\mu,A)}(y)=\big((A^{\top})^{-1}\big)^{\top}\nabla_{x}\log f(x)\big|_{x=(A^{\top})^{-1}(y-\mu)}=A^{-1}\nabla\log f\big((A^{\top})^{-1}(y-\mu)\big).$
Evaluating at $y_{i}=A^{\top}x_{\sigma(i)}+\mu$ gives
$\nabla_{y}\log f^{(\mu,A)}(y_{i})=A^{-1}\nabla\log f(x_{\sigma(i)}).$
Thus,
$S(Y)=\begin{pmatrix}\nabla_{y}\log f^{(\mu,A)}(y_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla_{y}\log f^{(\mu,A)}(y_{n})\end{pmatrix}=\begin{pmatrix}A^{-1}\nabla\log f(x_{\sigma(1)})\\[-2.0pt] \vdots\\[-2.0pt] A^{-1}\nabla\log f(x_{\sigma(n)})\end{pmatrix}=P\,A^{-1}\,S(X),$
which proves (15). ∎
Proposition A.2 (Attention can compute KDE).
Let $X\in\mathbb{R}^{n\times d}$ be the normalized data matrix with $\|x_{i}\|=1$ for all $i$. Consider a single self-attention head
$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(QK^{\top})\,V$
with
$Q=K=\tfrac{1}{h}X,\qquad V=I_{n},$
for some bandwidth parameter $h>0$. Then
$\mathrm{Attention}(Q,K,V)]_{ij}=\frac{\exp\!\left(-\frac{\|x_{i}-x_{j}\|^{2}}{2h^{2}}\right)}{\sum_{k=1}^{n}\exp\!\left(-\frac{\|x_{i}-x_{k}\|^{2}}{2h^{2}}\right)},$
i.e. the attention matrix coincides with the normalized Gaussian KDE kernel of bandwidth $h$.
Proof.
With $Q=K=X/h$, we have $QK^{\top}=XX^{\top}/h^{2}$, so
$\mathrm{Softmax}(QK^{\top})_{ij}=\frac{\exp(x_{i}^{\top}x_{j}/h^{2})}{\sum_{k}\exp(x_{i}^{\top}x_{k}/h^{2})}.$
Because all $\|x_{i}\|=1$,
$x_{i}^{\top}x_{j}=1-\tfrac{1}{2}\|x_{i}-x_{j}\|^{2}.$
Substituting and simplifying gives
$\exp(x_{i}^{\top}x_{j}/h^{2})=e^{1/h^{2}}\exp(-\|x_{i}-x_{j}\|^{2}/2h^{2}),$
and the common factor $e^{1/h^{2}}$ cancels in the row-wise softmax normalization. Hence
$\mathrm{Softmax}(QK^{\top})_{ij}=\frac{\exp(-\|x_{i}-x_{j}\|^{2}/2h^{2})}{\sum_{k}\exp(-\|x_{i}-x_{k}\|^{2}/2h^{2})}.$
Since $V=I_{n}$, multiplying by $V$ leaves this matrix unchanged, and the claim follows. ∎
Appendix B Attention visualization
Figure 12: We visualize attention of the eight individual heads of layer 0 as a heatmap. Yellow color denotes higher attention weight and blue is lower. We choose the particle ordering so that nearby particles are close in the ordering. Heads 0, 2, and 5 specialize on nearby points, heads 1 and 6 specialize on far-away points, whereas heads 3, 4, and 7 attend in specific directions.
Figure 13: We visualize attention of the eight individual heads of layer 0 as a scatter plot. We choose a random query point, marked with a red x. Yellow color denotes higher attention weight and blue is lower. Heads 0, 2, and 5 specialize on nearby points, head 1 specializes on far-away points, whereas heads 3, 4, 6, and 7 attend in specific directions.
To interpret the outputs of the Transformer model we visualize the attention of the eight individual heads in layer 0 in two different ways. First, in Figure 12 we show the full attention matrices as heatmaps. We choose the particle ordering so that nearby particles are close in the ordering. Second, in Figure 13 we choose a random query point, marked with a red x, and color other points according to the strength of the query point’s attention. We clearly observe an emergent behavior – heads specialize in different tasks. Head 1 specialized to look at far-away points, whereas Heads 0, 2, and 5 specialized to look at close- and mid-range interactions. Finally, Heads 3, 4, 6, and 7 specialized to look in specific directions. This emergent behavior further links the multi-head Transformer to kernel-based methods, but with multi-scale learned kernels.
Appendix C Comparison with the Score Matching Loss
The Hyvärinen score matching loss (Hyvärinen, 2005) estimates the data score without access to the true density by minimizing
$\mathcal{L}_{\mathrm{SM}}(\theta)=\mathbb{E}_{x\sim f(x)}\!\left[\|\,s_{\theta}(x)\,\|^{2}+2\,\nabla_{x}\!\cdot s_{\theta}(x)\right],$
which corresponds to the Fisher divergence $\mathbb{E}_{f}[\|s_{\theta}(x)-\nabla_{x}\log f(x)\|^{2}]$ up to a constant.
To avoid explicitly computing the divergence term, one can apply a finite-difference denoising trick, replacing $\nabla\!\cdot s_{\theta}(x)$ with a stochastic estimator
$\nabla\!\cdot s_{\theta}(x)\;\approx\;\mathbb{E}_{z\sim\mathcal{N}(0,I)}\!\left[\frac{z^{\top}\big(s_{\theta}(x+\alpha z)-s_{\theta}(x-\alpha z)\big)}{2\alpha}\right],$
yielding a practical divergence-free approximation used in many implementations of score matching.
In Figure 14 we compare the performance of the score matching loss compared to the proposed Transformer model. The Transformer model is superior in two respects. First, it does not need to be retrained on new samples. Second, the score matching loss is easy to under- or over-fit, as demonstrated in the figure by training for 10, 100 and 1000 steps.
Figure 14: Comparison of sliced score matching and our transformer model. The transformer model can be used without retraining and does not suffer from overfitting. We plot the negated score for ease of visualization.
Appendix D Runtime comparison
Figure 15: Runtime comparison between KDE and the Transformer model in 2 and 10 dimensions. Both are $O(n^{2})$ asymptotically, but empirically the Transformer scales better and has improved memory efficiency. KDE encounters an OOM error at $n=2^{15}$.
While KDE is known to fail in moderate and high dimensions, it is thought to be very fast. We compare the wall-clock runtime of KDE score estimation as in equation (10), and the Transformer model we propose. The Transformer model contains $800,000$ learned parameters. We use both KDE and the Transformer model on a single L40S GPU with 48Gb of memory. Both methods are $O(n^{2})$ asymptotically due to pairwise computations (kernel evaluations for KDE, attention for Transformers). However, in practice the observed scaling differs: Figure 15 shows that while KDE is faster for small sample size $n\leq 2048$, it becomes slower after this point, especially in higher dimension. Additionally, the naive KDE implementation runs out of GPU memory after $n=32,768$, while the Transformer benefits from the highly optimized attention kernels in modern deep learning frameworks, which are difficult to match with custom KDE implementations in practice.